OpenStudy (anonymous):

Find critical numbers of the function g(x)=x√4-x

OpenStudy (anonymous):

g(x)=x(4-x)^1/2 g'(x)=(4-x)^1/2+[1/2(4-x)^-1/2] what do i do next

hartnn (hartnn):

you mean critical point , right ? so, Take the derivative of function to find g'(x) and equate it to 0 g'(x) = 0 then the values of x are the critical poitns

hartnn (hartnn):

the 2nd term of your derivative is incorrect, maybe check again ?

OpenStudy (anonymous):

\[g(x)=x(4-x)^\frac{ 1 }{ 2 }\] \[g'(x)=(4-x)^\frac{ 1 }{ 2 }+ \frac{ 1 }{ 2 }x(4-x)^-\frac{ 1 }{ 2 }\]

hartnn (hartnn):

now thats correct :)

hartnn (hartnn):

wait did you miss the negative sign for the 2nd term ?

OpenStudy (anonymous):

\[g′(x)=(4−x)^\frac{ 1 }{ 2}+[\frac{ 1 }{ 2 } x(4−x)^-\frac{ 1 }{ 2 }]\]

hartnn (hartnn):

because derivative of 4-x is -1

hartnn (hartnn):

won't you get \(g′(x)=(4−x)^\dfrac{ 1 }{ 2}-[\dfrac{ 1 }{ 2 } x(4−x)^{-\dfrac{ 1 }{ 2 }}]\) ??

hartnn (hartnn):

then equating g'(x) = 0 gives \(g′(x)=(4−x)^\dfrac{ 1 }{ 2}-[\dfrac{ 1 }{ 2 } x(4−x)^{-\dfrac{ 1 }{ 2 }}]=0 \\or, (4-x) = x/2\)

OpenStudy (anonymous):

It is only a negative when you are doing the quotient rule. This is the product rule that we are using.

hartnn (hartnn):

ys, and the chain rule too! heard of it ?

OpenStudy (anonymous):

Oh yea

hartnn (hartnn):

d/dx (sqrt (4-x)) = 1/2sqrt (4-x) d/dx (4-x) = - 1/2 sqrt(4-x)

hartnn (hartnn):

so can you continue now ?

hartnn (hartnn):

yes/no/idk ?

OpenStudy (anonymous):

@hartnn is correct, the second term will be negative due to the chain rule.